# Quantum State Fidelity for Functional Neural Network Construction

### v1

@@ Abstract
Neuroscientists face challenges in analyzing high-dimensional neural recording data of dense functional networks. Without ground-truth reference data, finding the best algorithm for recovering neurologically relevant networks remains an open question. We implemented hybrid quantum algorithms to construct functional networks and compared them with the results of documented classical techniques. We demonstrated that our quantum state fidelity can provide a competitive alternative to classical metrics by revealing distinct functional networks. Our results suggest that quantum computing offers a viable and potentially advantageous alternative for data-driven modeling in neuroscience, underscoring its broader applicability in high-dimensional graph inference and complex system analysis.
@@ V Discussion
Our results on the IBM Kingston QPU demonstrate that quantum state fidelity circuits can be run efficiently on current NISQ hardware and point toward future scalability of our method on larger neural connectomes and quantum embeddings. Notably, each state fidelity calculation can be computed computed independently for each pair of neurons, which enables parallel execution across multiple devices and hardware platforms. Our method may also be robust to noise, as distance metrics computed with DDD were not significantly different from those without DDD. However, as quantum metrics on hardware were more correlated with classical metrics than quantum metrics on simulators, further improvements in quantum hardware are needed to larger network analyses.

### v2

@@ Abstract
Neuroscientists face challenges in analyzing high-dimensional neural recording data of dense functional networks. Without ground-truth reference data, finding the best algorithm for recovering neurologically relevant networks remains an open question. We implemented hybrid quantum algorithms to construct functional networks and compared them with the results of documented classical techniques. We demonstrated that our quantum state fidelity methods can provide competitive alternatives to classical metrics by revealing distinct functional networks. Our results suggest that quantum computing offers a viable and potentially advantageous alternative for data-driven modeling in neuroscience, underscoring its broader applicability in high-dimensional graph inference and complex system analysis.
@@ V Discussion
Our results on the IBM Kingston QPU demonstrate that quantum state fidelity circuits can be run efficiently on current NISQ hardware and point toward future scalability of our method on larger neural connectomes and quantum embeddings. Notably, each state fidelity calculation can be computed computed independently for each pair of neurons, which enables parallel execution across multiple devices and hardware platforms. Our method may also be robust to noise, as distance metrics computed with DDD were not significantly different from those without DDD. However, as quantum metrics were more correlated with classical metrics on hardware than on simulators, further improvements in quantum hardware are needed before achieving clinical relevance.